OpenStudy (anonymous):

1/2:2/3:3/4=? please let me know the answer soon it is urgent

OpenStudy (curry):

are u multipling

OpenStudy (curry):

or do ujust want to know the decimal value

OpenStudy (anonymous):

it says 6:8:9 i want details

OpenStudy (anonymous):

1/2 :2/3 :3/4 = 6 : 8 : 9.

OpenStudy (anonymous):

how do we get 1/2:2/3:3/4=6:8:9

jhonyy9 (jhonyy9):

- so if you see the line of nominator and denominator you get numbers what can being simplified - if you want this till n-1/n the result what you will get will be 1/n

jhonyy9 (jhonyy9):

- so if you have tought 1/2 multiplied 2/4 and ....

jhonyy9 (jhonyy9):

sorry with 2/3

OpenStudy (anonymous):

If Rs. 782 be divided into three parts, proportional to 1/2 :2/3 :3/4 , then the first part is:

OpenStudy (anonymous):

jhonyy9 please reply soon

jhonyy9 (jhonyy9):

what do you wrote (1/2)/(2/3)/(3/4)

jhonyy9 (jhonyy9):

hence is correct ?

jhonyy9 (jhonyy9):

so if yes this will be 1/2 x 3/2 =3/4 and this 3/4/3/4=3/4 x 4/3 = 1